# Finite Groups of Symplectic Automorphisms of Supersingular K3 surfaces in Odd Characteristics

### v2

@@ 3.3. A Lattice-Theoretic Interpretation of Symplectic Automorphism Groups
up to automorphisms of $\operatorname{II}_{9,41}$. Here the indices $9,41$ are simply chosen large enough to fulfill Nikulin’s criterion. We denote by $\mathbb{D}_{p,\sigma}$ the orthogonal complement of $N_{p,\sigma}$ in $\operatorname{II}_{9,41}$. Then $q_{\small{\mathbb{D}_{p,\sigma}}}=-q_{N_{p,\sigma}}$ and $\operatorname{sgn}(\mathbb{D}_{p,\sigma})=(8,20)$. Such a lattice $\mathbb{D}_{p,\sigma}$ is unique up to isomorphism and the orthogonal complement of any primitive embedding of $\mathbb{D}_{p,\sigma}$ in $\operatorname{II}_{9,41}$ is isomorphic to $N_{p,\sigma}$. We call $\mathbb{D}_{p,\sigma}$ the auxiliary lattice. The next proposition will be used to prove Theorem 3.8.
@@
The following theorem is a criterion of Leech pairs arising from symplectic actions of finite groups on supersingular K3 surfaces:
Theorem  3.8.
Let $(G,S)$ be a Leech pair with $\operatorname{rank}(S)\leq 21$. The following are equivalent:
  1. (1)
    for each supersingular K3 surface $X/k$ with Artin invariant $\sigma$, there is an embedding $G\hookrightarrow\operatorname{Aut}^{s}(X)$, with $(G,S_{G}(X))\cong(G,S)$;
  2. (2)
    there exists a supersingular K3 surface $X/k$ with Artin invariant $\sigma$, such that there is an embedding $G\hookrightarrow\operatorname{Aut}^{s}(X)$, with $(G,S_{G}(X))\cong(G,S)$;
  3. (3)
    there exists a primitive embedding $S\hookrightarrow N_{p,\sigma}$.
Proof.
@@ 3.4. Superspecial K3 Surfaces
In this subsection, we discuss the case of superspecial K3 surfaces, i.e., $\sigma=1$. Maximal finite subgroups of $\operatorname{Aut}^{s}(X)$ for superspecial K3 surfaces over an algebraically closed field of any characteristic (including $2$) are obtained recently by Ohashi–Schütt [OS24]. For further use in this paper, we give a larger list (Table LABEL:table:_sublist_of_HM) consisting of all Höhn–Mason pairs $(G,S)$ which admit primitive embedding $S\hookrightarrow N_{p,1}$ for odd prime $p$. Here Höhn–Mason pair is defined as follows:
Definition  3.9.
A Höhn–Mason pair $(G,S)$ is a Leech pair with $G\leq\operatorname{Co}_{0}$ and $S=S_{G}(\mathbb{L})$, such that $G$ is maximal in the sense that $G$ contains all automorphisms of $S$ acting trivially on $A_{S}$, i.e., $G$ has no nontrivial enlargement.
Such kind of pairs have been classified by Höhn–Mason [HM16] and there are 290 of them. Now we relate finite groups of symplectic automorphisms of a superspecial K3 surface with Höhn–Mason pairs. Recall that for any Leech pair $(G,S)$, we let $\widetilde{G}$ be the group of all automorphisms of $S$ acting trivially on $A_{S}$. The enlargement $(\widetilde{G},S)$ is still a Leech pair.
Proposition  3.10.
Suppose $X$ is a superspecial K3 surface and $G\leq\operatorname{Aut}^{s}(X)$ is finite. Then the pair $(G,S_{G}(X))$ is a Leech pair and it can be enlarged to a Höhn–Mason pair $(\widetilde{G},S_{G}(X))$.
@@
By Theorem 3.8, a Leech pair $(G,S)$ arises from a superspecial K3 surface in characteristic $p$ if and only if there exists a primitive embedding $S\hookrightarrow N_{p,1}$.
Theorem  3.11.
@@
Remark  3.12.
It is fine to replace the auxiliary lattice $\mathbb{D}_{p,1}$ by a negative even lattice of rank $4$ with same discriminant form.
Remark  3.13.
The results in Table LABEL:table:_sublist_of_HM with $\operatorname{rank}(S)=20,21$ have been obtained by Ohashi–Schütt [OS24], which is sufficient for their classification. The cases with $\operatorname{rank}(S)\leq 19$ will be useful to understand the size of finite groups of symplectic automorphisms of supersingular K3 surfaces with arbitrary Artin invariants, see Proposition 3.15 and §6. In the proof of Theorem 3.11, our description of the algorithm is brief. We refer to [OS24] for more details about the calculation.
Example  3.14.
@@ 3.5. Artin Invariants $\sigma\geq 2$
Proposition  3.15.
The enlargement $(\widetilde{G}_{L},S_{G_{L}}(N_{p,1}))$ of $(G_{L},S_{G_{L}}(N_{p,1}))$ is a Höhn–Mason pair belonging to Table LABEL:table:_sublist_of_HM.
@@ 4. The Tame Case
In this section we give the proof of Theorem 1.1.
4.1. Proof of Theorem 1.1 for Supersingular K3 Surfaces
The next lemma will be used in the proof of Theorem 1.1 for supersingular K3 surfaces.
@@
By Lemma 3.7, the pair $(G,S_{G}(X))$ is a Leech pair. We have $\operatorname{rank}(N_{p,1})+\ell_{q}(A_{N_{p,1}})=22$ if $q\neq p$. By Lemma 2.10 we have $\operatorname{rank}(S_{G}(N_{p,1}))+\ell_{q}(A_{S_{G}(N_{p,1})})\leq 22$. It is clear that $\operatorname{rank}(S_{G}(N_{p,1}))+\ell_{p}(A_{S_{G}(N_{p,1})})\leq 22$ since $\ell_{p}(A_{S_{G}(N_{p,1})})=0$. Therefore, $\operatorname{rank}(S_{G}(N_{p,1}))+\ell(A_{S_{G}(N_{p,1})})\leq 22$. From Lemma 4.1 and Proposition 2.9, the Leech pair $(G,S_{G}(X))$ is isomorphic to a subpair of $(\operatorname{Co}_{0},\mathbb{L})$. Since $S_{G}(X)$ admits a primitive embedding into $N_{p,1}$, the enlargement of $(G,S_{G}(X))$ is a Höhn–Mason pair belonging to Table LABEL:table:_sublist_of_HM.
We have a saturation $S_{G}(X)\oplus\operatorname{NS}(X)^{G}\subset\operatorname{NS}(X)$. By Lemma 4.1, $\ell_{p}(A_{S_{G}(X)})=0$, hence $\ell_{p}(\operatorname{NS}(X)^{G})=\ell_{p}(\operatorname{NS}(X))=2\sigma$. This implies that $\operatorname{rank}(\operatorname{NS}(X)^{G})\geq 2\sigma$. We conclude that $\operatorname{rank}(S_{G}(X))\leq 22-2\sigma$. Now the proof of Theorem 1.1 in the supersingular case is complete.
4.2. Proof of Theorem 1.1 for K3 Surfaces of Finite Height
In this subsection, we prove Theorem 1.1 for K3 surfaces of finite height $h$. We mainly use Jang [Jan17] to turn this into a problem over characteristic 0. For more on K3 surfaces of finite height, see [AM77] or [Lie16, §6] and references therein.
@@
Next we show that $g$ acts trivially on $A_{\operatorname{NS}(\mathscr{X}_{\overline{K}})}$. Since such a conclusion is true over $\mathbb{C}$, we can use the ”Lefschetz principle” here to conclude over an algebraically closed field of characteristic 0. To be more precise, there exists a finitely generated field $E$ over $\mathbb{Q}$ contained in $\bar{K}$ and a K3 surfaces $X^{\prime}$ defined over $E$ with $X^{\prime}\times_{E}\overline{K}\cong\mathscr{X}_{\overline{K}}$ and $g$ acts on $X^{\prime}$ symplectically and of finite order. Since $E$ is finitely generated over $\mathbb{Q}$, there is also a field embedding $E\hookrightarrow\mathbb{C}$. We denote by $X^{\prime}_{\mathbb{C}}=X^{\prime}\times_{E}\mathbb{C}$. By [Huy16a, §17, Lemma 2.2], we have $\operatorname{NS}(\mathscr{X}_{\overline{K}})\cong\operatorname{NS}(X^{\prime}_{\overline{E}})\cong\operatorname{NS}(X^{\prime}_{\mathbb{C}})$, and the action of $g$ on $X^{\prime}_{\mathbb{C}}$ is symplectic. We then know that the induced action of $g$ on $H^{2}(X^{\prime}_{\mathbb{C}},\mathbb{Z})$ fixes all elements in the transcendental lattice $T(X^{\prime}_{\mathbb{C}})$. Therefore, the induced action of $g$ on $A_{\operatorname{NS}(X)}\cong A_{\operatorname{NS}(X^{\prime}_{\mathbb{C}})}\cong A_{T(X^{\prime}_{\mathbb{C}})}$ is trivial.
@@
Finally, since $\gcd(|G|,p)=1$, we know $\gcd(|A_{S_{G}(X)}|,p)=1$, similar as Lemma 4.1. We hence finish the proof of Theorem 1.1 for K3 surfaces of finite height.
@@ 5. $p$-Root Pairs
In the following, we will first discuss irreducible $p$-root pairs and then general $p$-root pairs. In particular, in the last subsection, we prove the second statement in Theorem 1.2.
@@ 5.1. $p$-Root Pairs of Type $A$
Remark  5.7.
This may not hold in general. See Lemma 5.14.
@@ 5.3. $p$-Root Pair of Type $D_{4}$
We have that $\operatorname{Aut}(D_{4})=S_{3}\ltimes W(D_{4})$. Consider the outer automorphisms: $x:\alpha_{3}\mapsto\alpha_{4}\mapsto\alpha_{1}\mapsto\alpha_{3},\alpha_{2}\mapsto\alpha_{2}$, $y:v_{4}\rightarrow-v_{4}$, i.e., $y:\alpha_{3}\mapsto\alpha_{4}$. Then the outer automorphism group $S_{3}$ is generated by $x,y$.
@@
If $\lambda$ is a root, then two of $\{(2a+b),(-a-b),2b,(-a)\}$ has to be 00 in $\mathbb{Z}/3\mathbb{Z}$ which is impossible. ∎
@@
For $R=E_{6}$, We first show that $H\cap W(E_{6})\neq\{id\}$. Otherwise $H\cap W(E_{6})=\{id\}$, then $|H|=2$, which implies $S_{H}(E_{6})$ is rootless by Lemma 5.15. So $R^{H}$ is nonzero, contradiction. By Proposition 5.17, $|H\cap W(E_{6})|$ is a power of $p$. Hence $p\big{|}|W(E_{6})|$, $p=3$ or $5$.
@@ 6. Application of $p$-Root Pairs
And by Lemma 3.16, $(G_{R},R)$ is a $p$-root pair. In this section, we apply our results on $p$-root pairs to show that when $p\geq 13$, we always have $(p,|G|)=1$, which recovers Dolgachev–Keum [DK09, Theorem 2.1] for supersingular K3. And then we calculate the upper bounds of the $p$-part of $G$ for $p=3,5,7,11$.
@@ 6.1. New Proof for Dolgachev–Keum’s Result
The following theorem is obtained by Dolgachev–Keum [DK09, Theorem 2.1] through a geometric argument using fixed loci and fibrations. For supersingular cases, we provide a lattice theoretic proof.
@@ 6.2. Upper Bounds of Wild Degrees
For $p=11,7$, by Proposition 5.11, 5.13 and 5.19, all components of $R$ are of type A. Since $\operatorname{rank}R\leq 22$, by Lemma 5.22 and Proposition 5.5, each component of $R$ is isomorphic to $A_{p-1}$.
For $p=5$, by Proposition 5.11, 5.13, there is no irreducible component of type $D$. Then the conclusion follows from Lemma 5.22, Proposition 5.5, 5.21.
For $p=3$, by Proposition 5.11, Lemma 5.14, $D_{4}$ is the only possibility of type D. The remaining conclusion follows from Lemma 5.18, 5.22 and Proposition 5.5. ∎
@@ Appendix A Höhn–Mason Pairs Admitting Primitive Embedding into $N_{p,1}$
Combining with Kondō [Kon06], in the case $(G,q_{S})\cong([2^{4}3].S_{5},8_{1}^{+1}3^{-1}5^{-1})$, we actually have $G\cong 2^{4}:(3\times S_{5})$.

### v3

@@ 3.3. A Lattice-Theoretic Interpretation of Symplectic Automorphism Groups
up to automorphisms of $\operatorname{II}_{9,41}$. Here the indices $9,41$ are simply chosen large enough to fulfill Nikulin’s criterion. We denote by $\mathbb{D}_{p,\sigma}$ the orthogonal complement of $N_{p,\sigma}$ in $\operatorname{II}_{9,41}$. Then $q_{\small{\mathbb{D}_{p,\sigma}}}=-q_{N_{p,\sigma}}$ and $\operatorname{sgn}(\mathbb{D}_{p,\sigma})=(8,20)$. Such a lattice $\mathbb{D}_{p,\sigma}$ is unique up to isomorphism and the orthogonal complement of any primitive embedding of $\mathbb{D}_{p,\sigma}$ in $\operatorname{II}_{9,41}$ is isomorphic to $N_{p,\sigma}$. We call $\mathbb{D}_{p,\sigma}$ the auxiliary lattice. The next proposition will be used to prove Proposition 1.1.
@@
We now prove Proposition 1.1 which is a criterion of Leech pairs arising from symplectic actions of finite groups on supersingular K3 surfaces.
Proof of Proposition 1.1.
@@ 3.4. Superspecial K3 Surfaces
In this subsection, we discuss the case of superspecial K3 surfaces, i.e., $\sigma=1$. Maximal finite subgroups of $\operatorname{Aut}^{s}(X)$ for superspecial K3 surfaces over an algebraically closed field of any characteristic (including $2$) are obtained recently by Ohashi–Schütt [OS24]. Together with [DK09a], they complete the classification of finite tame symplectic actions on all K3 surfaces in positive characteristic. We define Höhn–Mason pair as follows:
Definition  3.8.
A Höhn–Mason pair $(G,S)$ is a Leech pair with $G\leq\operatorname{Co}_{0}$ and $S=S_{G}(\mathbb{L})$, such that $G$ is maximal in the sense that $G$ contains all automorphisms of $S$ acting trivially on $A_{S}$, i.e., $G$ has no nontrivial saturation.
Such kind of pairs have been classified by Höhn–Mason [HM16] and there are 290 of them. Now we relate finite groups of symplectic automorphisms of a superspecial K3 surface with Höhn–Mason pairs. Recall that for any Leech pair $(G,S)$, we let $\widetilde{G}$ be the group of all automorphisms of $S$ acting trivially on $A_{S}$. The saturation $(\widetilde{G},S)$ is still a Leech pair.
Proposition  3.9.
Suppose $X$ is a superspecial K3 surface and $G\leq\operatorname{Aut}^{s}(X)$ is finite. Then the pair $(G,S_{G}(X))$ is a Leech pair and it can be saturated to a Höhn–Mason pair $(\widetilde{G},S_{G}(X))$.
@@
By Proposition 1.1, a Leech pair $(G,S)$ arises from a superspecial K3 surface in characteristic $p$ if and only if there exists a primitive embedding $S\hookrightarrow N_{p,1}$.
Proposition  3.10.
@@
The calculation method is similar as the proof of [LZ22, Theorem 4.14], where Laza and the second author classified symplectic automorphism groups of smooth cubic fourfolds.
@@
Example  3.11.
@@ 3.5. Artin Invariants $\sigma\geq 2$
Proposition  3.12.
The saturation $(\widetilde{G}_{L},S_{G_{L}}(N_{p,1}))$ of $(G_{L},S_{G_{L}}(N_{p,1}))$ is a Höhn–Mason pair belonging to Table LABEL:table:_sublist_of_HM.
@@ 4. The Tame Case
In this section we give the proof of Proposition 1.2 and Proposition 1.3.
4.1. Proof of Proposition 1.2: Supersingular K3 Surfaces
The next lemma will be useful.
@@
By Lemma 3.7, the pair $(G,S_{G}(X))$ is a Leech pair. We have $\operatorname{rank}(N_{p,1})+\ell_{q}(A_{N_{p,1}})=22$ if $q\neq p$. By Lemma 2.10 we have $\operatorname{rank}(S_{G}(N_{p,1}))+\ell_{q}(A_{S_{G}(N_{p,1})})\leq 22$. It is clear that $\operatorname{rank}(S_{G}(N_{p,1}))+\ell_{p}(A_{S_{G}(N_{p,1})})\leq 22$ since $\ell_{p}(A_{S_{G}(N_{p,1})})=0$. Therefore, $\operatorname{rank}(S_{G}(N_{p,1}))+\ell(A_{S_{G}(N_{p,1})})\leq 22$. From Lemma 4.1 and Proposition 2.9, the Leech pair $(G,S_{G}(X))$ is isomorphic to a subpair of $(\operatorname{Co}_{0},\mathbb{L})$. Since $S_{G}(X)$ admits a primitive embedding into $N_{p,1}$, the saturation of $(G,S_{G}(X))$ is a Höhn–Mason pair belonging to Table LABEL:table:_sublist_of_HM.
We have a saturation $S_{G}(X)\oplus\operatorname{NS}(X)^{G}\subset\operatorname{NS}(X)$. By Lemma 4.1, $\ell_{p}(A_{S_{G}(X)})=0$, hence $\ell_{p}(\operatorname{NS}(X)^{G})=\ell_{p}(\operatorname{NS}(X))=2\sigma$. This implies that $\operatorname{rank}(\operatorname{NS}(X)^{G})\geq 2\sigma$. We conclude that $\operatorname{rank}(S_{G}(X))\leq 22-2\sigma$. Now the proof of Proposition 1.2 is complete.
4.2. Proof of Proposition 1.3: K3 Surfaces of Finite Height
We mainly use Jang [Jan17] to turn this into a problem over characteristic 0. For more on K3 surfaces of finite height, see [AM77] or [Lie16, §6] and references therein.
@@
Next we show that $g$ acts trivially on $A_{\operatorname{NS}(\mathscr{X}_{\overline{K}})}$. Since such a conclusion is true over $\mathbb{C}$, we can use the ”Lefschetz principle” here to conclude over an algebraically closed field of characteristic 0. To be more precise, there exists a finitely generated field $E$ over $\mathbb{Q}$ contained in $\bar{K}$ and a K3 surface $X^{\prime}$ defined over $E$ with $X^{\prime}\times_{E}\overline{K}\cong\mathscr{X}_{\overline{K}}$ and $g$ acts on $X^{\prime}$ symplectically and of finite order. Since $E$ is finitely generated over $\mathbb{Q}$, there is also a field embedding $E\hookrightarrow\mathbb{C}$. We denote by $X^{\prime}_{\mathbb{C}}=X^{\prime}\times_{E}\mathbb{C}$. By [Huy16, §17, Lemma 2.2], we have $\operatorname{NS}(\mathscr{X}_{\overline{K}})\cong\operatorname{NS}(X^{\prime}_{\overline{E}})\cong\operatorname{NS}(X^{\prime}_{\mathbb{C}})$, and the action of $g$ on $X^{\prime}_{\mathbb{C}}$ is symplectic. We then know that the induced action of $g$ on $H^{2}(X^{\prime}_{\mathbb{C}},\mathbb{Z})$ fixes all elements in the transcendental lattice $T(X^{\prime}_{\mathbb{C}})$. Therefore, the induced action of $g$ on $A_{\operatorname{NS}(X)}\cong A_{\operatorname{NS}(X^{\prime}_{\mathbb{C}})}\cong A_{T(X^{\prime}_{\mathbb{C}})}$ is trivial.
@@
Finally, since $\gcd(|G|,p)=1$, we know $\gcd(|A_{S_{G}(X)}|,p)=1$, similar as Lemma 4.1. We hence finish the proof of Proposition 1.3 for K3 surfaces of finite height.
@@ 5. $p$-Root Pairs
Lemma  5.2.
Proof.
In the following, we will first discuss irreducible $p$-root pairs and then general $p$-root pairs. In particular, in the last subsection, we prove the last statement in Theorem 1.5. Actually we here show that it holds more generally for pseudo $p$-root pairs .
@@ 5.1. $p$-Root Pairs of Type $A$
Remark  5.8.
This may not hold for other root systems. See Lemma 5.15.
@@ 5.3. $p$-Root Pair of Type $D_{4}$
We have that $\operatorname{Aut}(D_{4})=S_{3}\ltimes W(D_{4})$. Consider the outer automorphisms: $x:\alpha_{3}\mapsto\alpha_{4}\mapsto\alpha_{1}\mapsto\alpha_{3},\alpha_{2}\mapsto\alpha_{2}$, $y:v_{4}\rightarrow-v_{4}$, i.e., $y:\alpha_{3}\mapsto\alpha_{4}\mapsto\alpha_{3}$. Then the outer automorphism group $S_{3}$ is generated by $x,y$.
@@
If $\lambda$ is a root, then exactly two of $\{(2a+b),(-a-b),2b,(-a)\}$ has to be 00 in $\mathbb{Z}/3\mathbb{Z}$ which is impossible. ∎
@@
For $R=E_{6}$, we first show that $H\cap W(E_{6})\neq\{id\}$. Suppose that $H\cap W(E_{6})=\{id\}$, then $|H|=2$, which implies $S_{H}(E_{6})$ is rootless by Lemma 5.16. So $R^{H}$ is nonzero, contradiction. By Proposition 5.18, $|H\cap W(E_{6})|$ is a power of $p$. Hence $p\big{|}|W(E_{6})|$, $p=3$ or $5$.
@@ 6. Application of $p$-Root Pairs
And by Lemma 3.13, $(G_{R},R)$ is a $p$-root pair. In this section, we apply our results on $p$-root pairs to show that when $p\geq 13$, we always have $(p,|G|)=1$. This provides an alternative proof of [DK09a, Theorem 2.1] for supersingular K3. And then we calculate the upper bounds of the $p$-part of $G$ for $p=3,5,7,11$.
@@ 6.1. New Proof for Dolgachev–Keum’s Result
The following theorem is obtained by Dolgachev–Keum [DK09a, Theorem 2.1] through a geometric argument using fixed loci and fibrations. For supersingular cases, we provide a lattice theoretic proof.
@@ 6.2. Upper Bounds of Wild Degrees
For $p=11,7$, by Proposition 5.12, 5.14 and 5.20, all components of $R$ are of type A. Since $\operatorname{rank}R\leq 22$, by Lemma 5.2 and Proposition 5.6, each component of $R$ is isomorphic to $A_{p-1}$.
For $p=5$, by Proposition 5.12, 5.14, there is no irreducible component of type $D$. Then the conclusion follows from Lemma 5.2, Proposition 5.6, 5.22.
For $p=3$, by Proposition 5.12, Lemma 5.15, $D_{4}$ is the only possibility of type D. The remaining conclusion follows from Lemma 5.19, 5.2 and Proposition 5.6. ∎
@@
Upper bounds listed here do not depend on the Artin invariant $\sigma$.
@@ Appendix A Höhn–Mason Pairs Admitting Primitive Embedding into $N_{p,1}$
In the case $(G,q_{S})\cong([2^{4}3].S_{5},8_{1}^{+1}3^{-1}5^{-1})$, a more precise description of $G$ has been discussed in [OS24, Remark 7.2].